# Teacher-AI Collaboration for Curating and Customizing Lesson Plans in Low-Resource Schools

### v2

@@ Abstract.
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Computer-assisted instruction
@@ 2.2. AI For Supporting Teachers
AI is increasingly being adopted in educational settings to assist teachers with a range of tasks, including grading, administrative support, generating instructional content, and providing personalized feedback (Wang et al., 2024; OECD, 2024; Dorottya Demszky, 2025). These applications promise to reduce teacher workload, improve instructional quality, and support differentiated instruction. However, their effectiveness depends heavily on curricular alignment, contextual integration, and teacher trust (Viberg et al., 2024; Bitkina et al., 2020a). At the same time, Salehi et al. (2024) caution that introducing AI into knowledge work can reconfigure labor and deskill practitioners. This is particularly relevant for teachers who make extensive use of AI in their knowledge work.
@@
In response to these challenges, researchers have developed structured co-design frameworks to better align AI outputs with classroom needs. For instance, LessonPlanner embeds Gagné’s instructional principles into interactive UIs (Fan et al., 2024), while RAG frameworks refine LLM outputs through iterative feedback loops (Tan et al., 2025; Zheng et al., 2024; Mullins et al., 2024). Fine-tuned models and mega-prompting strategies that incorporate learner profiles, instructional goals, and teaching styles have also improved output fidelity (Karpouzis et al., 2024; Koenig, 2024). However, despite these advances, most research and design efforts to use AI to improve teachers’ skills and capacity remain primarily concentrated in Western, English-dominant contexts (Bucchiarone et al., 2024). In multilingual and low-resource classrooms, these systems often struggle because of lack of localized datasets, insufficient language support, and unique challenges faced by the teachers (Sebastian et al., 2024).
@@ 3. System Design and Implementation
Shiksha Copilot was developed to support teachers in low-resource Indian classrooms by enhancing lesson planning workflows through an LLM-based system with human-in-the-loop curation and bilingual support. The system was developed in close collaboration with an implementation partner that runs several initiatives to strengthen public education in Karnataka, India for over 22 years. Guru was piloted in Karnataka as part of the organization’s larger vision of making quality teaching and learning resources accessible to English- and Kannada-medium public school teachers. Aside from Guru, the interventions designed by the partner organization are implemented on the ground by mentors, who serve as conduits between the organization’s program team and public school teachers. These mentors train the teachers and provide ongoing support across all programs. The same approach was followed for Guru. The costs for lesson plan curation, tool deployment with teachers, and LLM API usage were funded by the partner organization through philanthropic grants.
@@ 3.1. System Design
The technical architecture of Guru is grounded in the core design principles outlined earlier and comprises three key components: Curriculum Ingestion, Translation, and LLM Orchestration.
@@ 4.2. Surveys
Survey Administration and Data Processing. The survey was administered via Qualtrics and shared with teachers through WhatsApp groups managed by mentors from the deployment team. Phone numbers served as identifiers to link pre- and post-intervention responses with lesson planning platform accounts. Participants were assured of confidentiality and informed that responses would not affect their professional standing. We received 601 complete responses for the pre-survey and 406 for the post-survey; analyses were conducted on the 406 matched responses. In terms of data cleaning, the missing values were imputed using miceforest (Mic, [n. d.]; Stekhoven and Bühlmann, 2012) and numerical outliers were winsorized (Sullivan et al., 2021) at the 95th percentile. All quantitative analyses were performed using Python-based statistical libraries. Composite scores for each construct were calculated as the mean of corresponding items. Full survey instruments are included in Appendix D.
@@ 5.4. Broader Impact on Teachers’ Work Practices
The adoption of Shiksha Copilot signaled a shift toward activity-based teaching. Teachers had been using the 5E lesson plan format even before the introduction of Shiksha Copilot, and they already recognized the value of classroom activities for fostering student engagement and deeper learning. However, they often struggled to design and implement such activities consistently due to limited time and a lack of readily available resources. Shiksha Copilot partially addressed this gap by providing easily accessible, contextually relevant activities tailored to each topic. These activities—such as visiting a gram panchayat office or collecting leaves with different venations—were grounded in the local environment and did not rely on digital infrastructure. Unlike device-dependent solutions such as Kahoot (Kah, [n. d.]) or Khanmigo (Kha, [n. d.]), the tool offered no-tech, practical ideas using commonly available materials, making them particularly well-suited for low-resource classrooms. One English teacher noted that the Clap and Tap activity—where students clap upon hearing consonant sounds and tap their heads for vowel sounds—was both engaging and required no materials. Similarly, physical activities like treasure hunts and relay races—where students search for concept-related clues around the school while competing with peers—supported embodied learning by combining cognitive and physical stimulation. Teachers emphasized that the availability of such ready-to-use activities reduced their lesson planning burden and motivated them to experiment with new instructional approaches. Teacher 15, a Social Science teacher, explained:
@@ 6.2. Leveraging the Communities of Practice
In India, this framework is particularly relevant given the scale and structure of the teaching profession. With nearly 10 million teachers across 1.47 million schools, the system is supported by over 600 District Institutes of Education and Training (DIETs), more than 6,000 Block Resource Centres (BRCs), and nearly 70,000 Cluster Resource Centres (CRCs) (Kaur, nd). These institutions form a hierarchical but collaborative support structure: DIETs coordinate district-level training, BRCs manage block-level programs, and CRCs facilitate localized mentoring and peer learning through regular cluster meetings (Wolfenden, 2015). However, despite this extensive infrastructure, professional development often remains uneven in quality and disconnected from teachers’ everyday realities (Dyer et al., 2004).

### v3

# Shiksha Copilot: Teacher-AI Collaboration for Curating and Customizing Lesson Plans in Low-Resource Schools

@@ Abstract.
copyright: ccjournal: PACMHCIjournalyear: 2026journalvolume: 10journalnumber: 2article: CSCW038publicationmonth: 4doi: 10.1145/3788074ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Computer-assisted instruction
@@ 2.2. AI For Supporting Teachers
AI is increasingly being adopted in educational settings to assist teachers with a range of tasks, including grading, administrative support, generating instructional content, and providing personalized feedback (Wang et al., 2024; OECD, 2024; Demszky et al., 2025). These applications promise to reduce teacher workload, improve instructional quality, and support differentiated instruction. However, their effectiveness depends heavily on curricular alignment, contextual integration, and teacher trust (Viberg et al., 2024; Bitkina et al., 2020a). At the same time, Salehi et al. (2024) caution that introducing AI into knowledge work can reconfigure labor and deskill practitioners. This is particularly relevant for teachers who make extensive use of AI in their knowledge work.
@@
In response to these challenges, researchers have developed structured co-design frameworks to better align AI outputs with classroom needs. For instance, LessonPlanner embeds Gagné’s instructional principles into interactive UIs (Fan et al., 2024), while RAG frameworks refine LLM outputs through iterative feedback loops (Tan et al., 2025; Zheng et al., 2024; Mullins et al., 2024). Fine-tuned models and mega-prompting strategies that incorporate learner profiles, instructional goals, and teaching styles have also improved output fidelity (Karpouzis et al., 2024; Koenig, 2024). However, despite these advances, most research and design efforts to use AI to improve teachers’ skills and capacity remain primarily concentrated in Western, English-dominant contexts (Bucchiarone and others, 2024). In multilingual and low-resource classrooms, these systems often struggle because of lack of localized datasets, insufficient language support, and unique challenges faced by the teachers (Sebastian et al., 2024).
@@ 3. System Design and Implementation
Shiksha Copilot was developed to support teachers in low-resource Indian classrooms by enhancing lesson planning workflows through an LLM-based system with human-in-the-loop curation and bilingual support. The system was developed in close collaboration with an implementation partner that runs several initiatives to strengthen public education in Karnataka, India for over 22 years. Shiksha Copilot was piloted in Karnataka as part of the organization’s larger vision of making quality teaching and learning resources accessible to English- and Kannada-medium public school teachers. Aside from Shiksha Copilot, the interventions designed by the partner organization are implemented on the ground by mentors, who serve as conduits between the organization’s program team and public school teachers. These mentors train the teachers and provide ongoing support across all programs. The same approach was followed for Shiksha Copilot. The costs for lesson plan curation, tool deployment with teachers, and LLM API usage were funded by the partner organization through philanthropic grants.
@@ 3.1. System Design
The technical architecture of Shiksha Copilot is grounded in the core design principles outlined earlier and comprises three key components: Curriculum Ingestion, Translation, and LLM Orchestration.
Refer to caption System flow diagram of Shiksha Copilot Detailed system architecture diagram of the Shiksha Copilot platform. On the left, teachers submit user queries and curators submit curation queries. Curriculum content is extracted from textbooks and stored as curriculum details and a curriculum content index in databases, alongside lesson plans and teacher profiles. Curators review and curate AI-generated lesson plans, which are stored back into the system. An orchestrator manages multiple AI agents, including retrieval-augmented generation agents, lesson plan generation and validation agents, web search and validation agents for Edu Chat, question generation agents for a question bank, and translation agents. These components work together to produce final Shiksha Copilot outputs such as lesson plans and lesson chat responses delivered back to teachers.
@@ 3.2. System Implementation
Refer to caption System workflow diagram System workflow diagram divided into a curation platform and a distribution platform. In the curation platform, a curator works with an LLM to curate learning objectives, develop an English lesson plan, and translate it into a Kannada lesson plan. These lesson plans are stored in a central lesson plans database. In the distribution platform, teachers use an LLM to customize lesson plans, resulting in individualized lesson plans for the classroom. Optional LLM-based features include lesson chat, educational chat, and question generation, which are available only to power users.
@@
Refer to caption User interface of the curator platform User interface screenshot of a lesson plan editing dashboard for a Kannada-medium Grade 6 mathematics lesson on fractions. At the top, metadata fields display medium, grade, subject, chapter, teaching method, lesson plan level, and last edited details. The main content area shows expandable lesson plan sections including subtopics, learning outcomes, and engage. The engage section is open and includes a Kannada text editing area on the left and an English reference version on the right. Instructions for saving edits and a prominent save changes button are visible. A navigation sidebar on the left provides access to lesson plan editing, history, and admin options.
@@
Refer to caption User interface screenshot of the teacher’s platform User interface screenshot of a teacher lesson plan view. A navigation sidebar on the left includes options such as dashboard, profile, content generation, chatbot, and question paper generation, along with lesson sections like learning outcomes, engage, explore, explain, elaborate, evaluate, and lesson summary. The main panel displays a lesson plan for algebraic expressions and identities, showing detailed learning outcomes in bullet form and an engage section with an introduction, key concepts, and a real-world gardening scenario. Chapter details and subtopics are visible in a side panel, and a chatbot button appears in the top-right corner.
@@ 4.2. Surveys
Survey Administration and Data Processing. The survey was administered via Qualtrics and shared with teachers through WhatsApp groups managed by mentors from the deployment team. Phone numbers served as identifiers to link pre- and post-intervention responses with lesson planning platform accounts. Participants were assured of confidentiality and informed that responses would not affect their professional standing. We received 601 complete responses for the pre-survey and 406 for the post-survey; analyses were conducted on the 406 matched responses. In terms of data cleaning, the missing values were imputed using miceforest (72; D. J. Stekhoven and P. Bühlmann (2012)) and numerical outliers were winsorized (Sullivan et al., 2021) at the 95th percentile. All quantitative analyses were performed using Python-based statistical libraries. Composite scores for each construct were calculated as the mean of corresponding items. Full survey instruments are included in Appendix D.
@@ 5.4. Broader Impact on Teachers’ Work Practices
The adoption of Shiksha Copilot signaled a shift toward activity-based teaching. Teachers had been using the 5E lesson plan format even before the introduction of Shiksha Copilot, and they already recognized the value of classroom activities for fostering student engagement and deeper learning. However, they often struggled to design and implement such activities consistently due to limited time and a lack of readily available resources. Shiksha Copilot partially addressed this gap by providing easily accessible, contextually relevant activities tailored to each topic. These activities—such as visiting a gram panchayat office or collecting leaves with different venations—were grounded in the local environment and did not rely on digital infrastructure. Unlike device-dependent solutions such as Kahoot (51) or Khanmigo (56), the tool offered no-tech, practical ideas using commonly available materials, making them particularly well-suited for low-resource classrooms. One English teacher noted that the Clap and Tap activity—where students clap upon hearing consonant sounds and tap their heads for vowel sounds—was both engaging and required no materials. Similarly, physical activities like treasure hunts and relay races—where students search for concept-related clues around the school while competing with peers—supported embodied learning by combining cognitive and physical stimulation. Teachers emphasized that the availability of such ready-to-use activities reduced their lesson planning burden and motivated them to experiment with new instructional approaches. Teacher 15, a Social Science teacher, explained:
@@ 6.2. Leveraging the Communities of Practice
In India, this framework is particularly relevant given the scale and structure of the teaching profession. With nearly 10 million teachers across 1.47 million schools, the system is supported by over 600 District Institutes of Education and Training (DIETs), more than 6,000 Block Resource Centres (BRCs), and nearly 70,000 Cluster Resource Centres (CRCs) (Kaur, n.d.). These institutions form a hierarchical but collaborative support structure: DIETs coordinate district-level training, BRCs manage block-level programs, and CRCs facilitate localized mentoring and peer learning through regular cluster meetings (Wolfenden, 2015). However, despite this extensive infrastructure, professional development often remains uneven in quality and disconnected from teachers’ everyday realities (Dyer et al., 2004).
@@ 7. Conclusion
Acknowledgements
We would like to thank the teachers who participated in the study for their time and thoughtful inputs. We extend our acknowledgments to Sikshana Foundation, especially Srinivas Mysore and Rijutha Raj, for their support in conducting the study. We thank Global Cornell and Cornell Global AI initiative for supporting this work. We also thank anonymous reviewers for their constructive and respectful feedback on our work.